A ball having mass of 100g fall down from a height of 1m and droped to the ground  and rebounds to the height of 40cm find impulsive force given time 0.1s 

Grade:12th pass

1 Answers

Khimraj
5 years ago
M = 100 gm
h = 1 m
velocity before touching the floor = u  m/s
   u² = 2 g h =  2 * 10 * 1 = 20
   u = - 20 m/s   downwards.
height of rebound = 0.4 m
velocity after touching the ground = v  m/s
 v² = 2 g h = 2* 10* 0.4 = 8 
 v = 2.5 m/s
Impulse is equal to the change in linear momentum during the impact/collision
     = m v - m u  = 100 /1000 kg * (2.5 - (-20)) = 2.25 kg-m/s  or N-sec